How do I show my work for y=x-4
4x+y=26

Answer by rfer(16322)

y=x-4
4x+y=26
4x+x-4=26
5x=30
x=30/5
x=6
------------
4*6+y=26
24+y=26
y=2
----------
y=x-4
2=6-4
2=2